I have problem with word problems and need some help.
Science and medication:
Ariana took 2 hrs longer to drive 360 miles on the first day of a vaction trip than to drive 144 miles the second day. If his averade driving rate was the same both days what was the driving time each day?

Answer by stanbon(75887)

Ariana took 2 hrs longer to drive 360 miles on the first day of a vaction trip than to drive 144 miles the second day. If his average driving rate was the same both days what was the driving time each day?
------------
1st day DATA:
time = x+2 hrs; distance = 360 miles; rate = d/t = 360/(x+2) mph
-------------
2nd day DATA::
time = x hrs; distance = 144 miles; rate = d/t = 144/x mph
---------------
EQUATION:
rate = rate
360/(x+2) = 144/x
360x = 144x + 288
216x = 288
x = 1 1/3 hrs (time on 2nd day)
x+2 = 3 1/3 hrs (time on 1st day)
============
Cheers,
Stan H.